Which should have a greater boiling point, a 1.0 m solution of acetic acid in water or a 0.75 m solution of NaCl in water. (Colligative Properties)
Answer #1

\DeltaTb   = i*Kb*m

CH3COOH (aq) ---------------> CH3COO^- (aq) + H^+ (aq)

i   = 2

Kb = 0.5120C/m

\DeltaTb   = i*Kb*m

           = 2*0.512*1

            = 1.0240C

The boiling point of CH3COOH in water = 100+ 1.024   = 101.0240C

NaCl(aq) ---------------> Na^+ (aq) + Cl^- (aq)

      i   = 2

\DeltaTb   = i*Kb*m

           = 2*0.512*0.75   = 0.7680C

The boiling point of NaCl in water = 100+ 0.768   = 100.7680C

1.0 m solution of acetic acid in water has higher boiling point than 1.0 m solution of acetic acid in water has higher boiling point than